Solve the system of equations by graphing. Then classify the system as consistent or inconsistent and as dependent or independent
x+y=16
x-y=-2

Answer by stanbon(75887)

Solve the system of equations by graphing. Then classify the system as consistent or inconsistent and as dependent or independent
x+y=16
x-y=-2
----
Rearrange:
y = -x+16
y = x+2
----
The two graphs have different slopes, so the
system is consistent and independent.
Cheers,
Stan H.